### v3

# Information Extraction from Clinical Notes: Are We Ready to Switch to Large Language Models?

Abstract
Backgrounds:
Information extraction (IE) is critical in clinical natural language processing (NLP). While large language models (LLMs) excel on generative tasks, their performance on extractive tasks remains debated.
Methods:
We investigated two fundamental clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). Using 1,588 clinical notes from four data sources including UT Physicians (UTP), MTSamples, MIMIC-III, and i2b2, we developed a comprehensive annotated corpus covering 4 main clinical entities (problems, tests, medications, and other treatments) and 16 modifiers (e.g., negation, certainty). We instruction-tuned LLaMA-2 and LLaMA-3 for both tasks, comparing their performance, generalizability, computational resources, and throughput to BiomedBERT.
Results:
LLaMA models consistently outperformed BiomedBERT across datasets and tasks. With sufficient training data (e.g., UTP), LLaMA showed marginal improvements (1% on NER, 1.5-3.7% on RE) over BiomedBERT; for limited training data (e.g., MTSamples, MIMIC-III), the improvements were greater. On the unseen i2b2 dataset, LLaMA-3-70B outperformed BiomedBERT by over 7% (F1) on NER and 4% (F1) on RE. However, LLaMA models required significantly more computing resources (i.e., memory and GPUs) and ran up to 28 times slower than BiomedBERT. To facilitate LLM use in clinical IE, we implemented ”Kiwi,” a clinical IE package featuring both BiomedBERT and LLaMA models and made it available at https://kiwi.clinicalnlp.org/.
Conclusion:
This study is among the first to develop and evaluate a comprehensive clinical IE system using open-source LLMs. Results indicate that LLaMA models outperform BiomedBERT for clinical NER and RE but with higher computational costs and lower throughputs. These findings highlight that choosing between LLMs and traditional deep learning methods for clinical IE applications should remain task-specific, taking into account both performance metrics and practical considerations such as available computing resources and the intended use case scenarios.
1 Introduction
The digital transformation of healthcare records into Electronic Health Records (EHRs) represents one of the most significant advancements in the modern healthcare system 1. This transformation has paved the way for approaches to advance patient care and clinical research 2. Among the data contained within EHRs, clinical notes have the most comprehensive and information-rich patient data 3. These narrative texts, written by healthcare professionals, capture a detailed picture of patient information, including diagnosis, treatments, patient outcomes, and much more 4. Despite their potential, the massive, complex and unstructured nature of clinical notes has largely presented significant challenges for their effective use 5.
Clinical information extraction (IE) aims to bridge the gap between unstructured text and structured data analysis 6. By automating this process using Natural Language Processing (NLP) techniques such as Named Entity Recognition (NER) and Relation Extraction (RE), clinical IE offers solutions to these challenges 7; 8. Over the past 20 years, most clinical NLP research has focused on clinical IE. For instance, 8 out of 10 challenges in the i2b2/n2c2 series from 2006 to 2019 have centered on clinical IE tasks 9; 10; 11; 12; 13; 14; 15; 16. BERT-based models have previously demonstrated state-of-the-art (SOTA) performance in clinical IE tasks 17; 18. However, the complexity and variability of natural language used by different healthcare professionals present substantial barriers to effective clinical IE 5. Systems developed using documents from a specific institution often do not generalize well when applied to clinical notes from other institutions or medical specialties 19.
Recently, large language models (LLMs) such as Generative Pre-trained Transformers (GPT) and the Large Language Model Meta AI (LLaMA) series 20; 21; 22 have shown phenomenal capabilities in generative NLP tasks, including text summarization, question answering, and machine translation 23; 24. However, the application of these models to traditional IE tasks has yielded mixed results, with outcomes varying across different studies and scenarios, sparking debates within the research community. For example, in zero-/few-shot settings, LLMs often do not outperform traditional fine-tuned BERT models 25; 26; 27. While some research indicates that LLMs can achieve comparable or slightly better performance than traditional models after fine-tuning, these findings are frequently based on datasets from similar text types, raising questions about the models’ generalizability across diverse text types 28; 29. Additionally, many studies focus on datasets with limited entity types or relation types 30; 31; 32. To the best of our knowledge, no comprehensive study has compared fine-tuned LLMs and BERT models in terms of performance and generalizability across clinical notes from multiple data sources, entity types, and relation types. Moreover, there is a lack of research examining the trade-offs between model performance, throughput, and energy consumption, which is particularly crucial given the varying computational resources available in different hospitals and research institutions.
In response to these challenges, we developed a comprehensive clinical IE corpus, covering major clinical entities, using 1,588 clinical notes from various institutions. This corpus includes 4 main entities and their 16 corresponding modifiers. Using our corpus, we instruction-tuned and evaluated LLaMA-2 and LLaMA-3 models for both NER and RE. We systematically compared these models with BiomedBERT in terms of performance, generalizability, throughput, and energy consumption to provide insights for users on selecting the right tool for clinical IE. To facilitate real-world application, we introduced “Kiwi”, a ready-to-use clinical IE pipeline that includes our instruction-tuned LLaMA-based and BiomedBERT-based models. An online demo is available at: https://kiwi.clinicalnlp.org/.
2 Methods
2.1 Task
Our study includes two clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). NER identifies entity types and boundaries in text 33, while RE identifies relationships between entity pairs 34. We instruction-tuned LLaMA-2 and LLaMA-3 models and compared them to fine-tuned BiomedBERT. We focused on extracting four main entities—Medical Problems, Tests, Drugs and other Treatments—along with 16 modifiers (e.g., Severity, Temporal, Dosage). In total, there are 21 possible relations between main entities and modifiers, as detailed in Table 1. Annotation methods and guidelines are available in the Supplementary File and GitHub repository.
UTP: UT Physicians; MTSamples: Transcribed Medical Transcription Sample Reports and Examples; MIMIC-III: Medical Information Mart for Intensive Care; i2b2: Informatics for Integrating Biology & the Bedside
Figure 1: (A) Overview of the study workflow, including entity and relation types in the dataset. (B) Examples of instructions for NER and RE, using a unified format with span tags to highlight all entities and modifiers. (C) The GUI of Kiwi online demo.
2.2 Dataset
We developed a comprehensive clinical IE corpus that is not limited to specific diseases or treatments. It consists of real-world and synthetic clinical notes collected cross-institutionally, including 1,342 clinical notes from UT Physicians (UTP), 146 from Transcribed Medical Transcription Sample Reports and Examples (MTSamples) 35, 50 from the Medical Information Mart for Intensive Care (MIMIC-III) 36, and 50 from 2010 Informatics for Integrating Biology & the Bedside (2010 i2b2) 9. Tables 1 presents the composition and statistics of the four datasets used in this study for NER and RE.
Train (Finetune) Test
UTP MTSamples MIMIC-III UTP MTSamples MIMIC-III i2b2
Documents 1,292 96 25 50 50 25 50
Main Entity Modifier
Problem 85,457 4,059 2,238 3,446 2,200 2,379 1,454
Negation 46,218 1,478 438 1,863 780 575 271
Temporal 2,737 101 156 114 65 181 73
Severity 1,544 114 91 53 38 131 37
Condition 296 17 42 18 20 74 1
Certainty 551 62 87 15 20 63 12
Subject 2,575 26 20 132 12 26 0
Body location 24,205 959 720 956 588 726 468
Course 2,971 220 163 114 109 184 93
Test 50,003 1,529 2,297 2,151 828 2,542 1,309
Negation 224 11 9 5 7 11 2
Temporal 24,027 115 2,130 1,055 30 3,057 474
Lab value 33,009 893 1,606 1,433 398 1,963 919
Reference range 2,287 0 11 44 0 5 0
Drug 13,648 747 1,047 568 466 1,021 681
Negation 321 20 7 6 8 15 3
Temporal 16,109 46 44 728 14 69 60
Strength 6,590 154 492 275 171 436 268
Dosage 5,352 61 258 231 46 248 99
Duration 514 27 77 21 11 27 32
Form 12,028 58 396 536 47 359 144
Frequency 6,440 209 521 262 219 434 320
Route 7,079 106 377 301 92 353 260
Other Treatment 8,208 584 542 329 388 510 405
Negation 321 12 16 6 5 19 6
Temporal 1,534 53 56 59 38 98 59
Table 1: Numbers of documents, entities, and combinations between main and modifier entities for train and test splits across all datasets.
2.3 Models
2.3.1 Large Language Model Meta AI (LLaMA)
We instruction-tuned LLaMA-2-chat and LLaMA-3-instruct models (7B, 13B, and 70B for LLaMA-2-chat; 8B and 70B for LLaMA-3-instruct) using Parameter-Efficient Fine Tuning (PEFT) to enhance computational efficiency 37. Specifically, we used 4-bit quantized models with QLoRA (LoRA settings: r=16, $\alpha=$64, dropout=0.05) and focused on tuning linear layers to reduce computational load 38. Fine-tuning parameters included a learning rate of 2e-4, 2 epochs, batch size of 4, and warmup ratio of 0.05. Model weights are available at huggingface.co/meta-llama 39, and we used the vLLM package with a temperature of 0 for reproducible output generation 40.
2.3.2 Bidirectional Encoder Representations from Transformers (BERT)
We evaluated BiomedBERT 41 for NER and RE tasks. BiomedBERT was selected over other clinical domain BERT models due to its better performance observed in our preliminary experiments. NER settings included a learning rate of 5e-5, 20 epochs, weight decay of 0.01, and batch size of 4. For RE, we used a learning rate of 1e-6, 20 epochs, and batch size of 64.
2.4 Instruction format
Our approach uses a unified format for NER and RE (see Figure 1(B)) as listed below: NER: LLaMA models are instructed to identify main entities in clinical note sentences, tagging them with <span class=’entity_type’>… </span>, where ”entity_type” defines the entity. RE: The main entity in the input text is tagged similarly, and LLaMA models identify related modifier entities using the same span format. This approach simplifies entity and modifier recognition in a single step. Full instructions are in Supplementary Table S1.
2.5 Evaluation
2.5.1 Performance and significance test
We evaluate RE and NER performance using Precision (P), Recall (R), and F1 scores under exact and relaxed match criteria. Exact matches require entity types, boundaries, and relations to align perfectly, while relaxed matches require only entity type and relation matches with overlapping boundaries. Significance tests assess improvements over BiomedBERT, calculating p-values via a one-sided Wilcoxon rank-sum test with bootstrapping (sample size equal to test size, 1000 repetitions, 95% confidence interval).
2.5.2 Cross-institution generalizability
We assess cross-institution generalizability using the i2b2 dataset, which contains clinical notes from Partners Healthcare and Beth Israel Deaconess Medical Center. This dataset’s diversity makes it ideal for testing model generalization to new institutions. Therefore, we excluded i2b2 from fine-tuning to use it solely for generalizability testing.
2.5.3 Computing costs
We assess each model’s computational costs by measuring GPU hours, memory usage, and energy consumption during fine-tuning and inference. GPU hours are calculated by multiplying the number of GPUs by the total processing time. Memory usage reflects the average GPU memory in GB per GPU. Energy consumption is estimated from average GPU power usage and converted to carbon emissions using a factor of 0.39 kg CO2)/kWh 42.
2.5.4 Throughput
We assess each model’s throughput during fine-tuning and inference. For fine-tuning, we measure the total time to train and calculate seconds per note by dividing the total time by the number of notes in the dataset. For inference, we determine the time per note using the test dataset with a batch size of 100 for both LLaMA and BiomedBERT models.
3 Results
3.1 Fine-tuning performance
Table 2 highlights the exact match performance of LLaMA-2, LLaMA-3, and BiomedBERT models, with full metrics in Supplementary Table S2. LLaMA models consistently outperformed BiomedBERT, particularly with limited fine-tuning data. With sufficient data (e.g., UTP dataset), LLaMA models show minor improvements of  1% on NER and 1.5-3.7% on RE. However, with less data (e.g., MTSamples, MIMIC-III), LLaMA models show greater performance improvement compared to BiomedBERT, with 3.7-4.5% higher F1 scores on NER and up to 12.5% higher on RE.
Task Model UTP MTSamples MIMIC-III i2b2 (Unseen)
NER LLaMA-2-7B 0.929 0.860 0.838 0.846
LLaMA-2-13B 0.932 0.868 0.847 0.853
LLaMA-2-70B 0.931 0.871 0.847 0.860
LLaMA-3-8B 0.929 0.869 0.843 0.852
LLaMA-3-70B 0.932 0.876 0.855 0.872
BiomedBERT 0.921 0.833 0.810 0.798
RE LLaMA-2-7B 0.916 0.785 0.823 0.730
LLaMA-2-13B 0.915 0.793 0.833 0.725
LLaMA-2-70B 0.918 0.795 0.850 0.736
LLaMA-3-8B 0.936 0.787 0.859 0.739
LLaMA-3-70B 0.937 0.795 0.858 0.744
BiomedBERT 0.898 0.670 0.808 0.703
Table 2: Exact F1 scores of LLaMA-2, LLaMA-3, and BiomedBERT on NER and RE across three datasets and cross-institution generalizability test results on unseen dataset (i2b2). All results of LLaMA models are statistically significant compared to BiomedBERT based on the one-sided Wilcoxon rank-sum test.
To explore the performance of LLaMA and BiomedBERT in low-resource settings, we conducted dataset-specific fine-tuning by training two models—LLaMA-3-70B and BiomedBERT—separately on the MTSamples and MIMIC-III datasets. Each model was fine-tuned exclusively on either MTSamples or MIMIC-III, then evaluated across all four test sets. Based on the results in Table 2, where LLaMA-3-70B demonstrated superior performance across datasets, we prioritized using this model for these experiments. Our findings show that LLaMA-3-70B consistently outperformed BiomedBERT across datasets in these low-resource scenarios, as shown in Supplementary Tables S3 and S4.
3.3 Computational resources and throughput
We compared the computational resources and throughput of BiomedBERT, LLaMA-2, and LLaMA-3 using NVIDIA A100 80GB GPUs (Tables 3 and 4). Fine-tuning was conducted with 1 GPU for BiomedBERT and the LLaMA. For inference, BiomedBERT and the 7B, 8B, and 13B LLaMA models used 1 GPU, whereas the 70B models required 2 GPUs. Results show notable trade-offs between performance and computational efficiency across model sizes.
3.3.1 Memory usage
Memory requirements increase substantially with model size. The 70B versions of LLaMA-2 and LLaMA-3 required 58-66 GB for training and 146 GB for inference, while BiomedBERT used only 21.3 GB and 65.7 GB, respectively. The 7B and 8B LLaMA models showed comparable memory efficiency to BiomedBERT, with training memory usage between 12.7-32.4 GB. Notably, during training, LLaMA models utilize Low-Rank Adaptation (LoRA), which reduces memory requirements compared to inference, where LoRA is not applied. This difference makes the memory needed for fine-tuning significantly lower than that required for inference. Additionally, these memory usages are heavily influenced by batch sizes; the current setup is optimized for best model performance.
3.3.2 Energy consumption and carbon emission
Energy consumption varied widely across models, with larger models requiring substantially more resources. LLaMA-70B models were the most energy-intensive during training, using 101-125 kWh and producing 39-49 kg of CO2). Smaller LLaMA models and BiomedBERT were more efficient, using 13-28 kWh and 5-11 kg of CO2). For inference, BiomedBERT was highly efficient at 0.03 kWh, while LLaMA models ranged from 0.14 kWh (LLaMA-3-8B) to 0.86 kWh (LLaMA-2-70B).
LLaMA2- 7B LLaMA2- 13B LLaMA2- 70B LLaMA3- 8B LLaMA3- 70B Biomed BERT
Memory Usage (GB)
Train 12.7 17.4 57.6 32.4 66.3 21.3
Test* 72.8 72.9 146.9 72.7 145.9 65.7
Total GPU Hours (hours)
Train 41.7 71.2 320.2 38.6 273.5 96.8
Test* 0.5 0.9 2.2 0.4 2.2 0.08
GPU Hours per Epoch (hours)
Train 20.85 35.6 160.1 19.3 136.75 4.84
Energy Consumption (kWh)
Train 15.22 27.98 125.21 13.16 101.21 15.7
Test* 0.19 0.36 0.86 0.14 0.81 0.03
Carbon Emissions (kg CO2)
Train 5.94 10.91 48.83 5.13 39.47 6.12
Test* 0.08 0.14 0.33 0.06 0.32 0.01
  • Training: 1413 documents; Test: 175 documents. *70B models need two A100 80GB GPUs for testing.
Table 3: Training and test times, and processing throughputs across models and datasets.
3.3.3 Training and Inference Throughput
Training and inference throughputs varied substantially across models and datasets as shown in table 4. LLaMA-3-8B had the fastest training time at 38.6 GPU hours, while BiomedBERT required 96.8 hours. BiomedBERT’s longer training time can be attributed to its need for 20 training epochs, compared to only 2 epochs for LLaMA models. This difference is reflected in GPU hours per epoch: BiomedBERT stands out with a particularly low requirement of 4.84 GPU hours per epoch, whereas LLaMA2-7B and LLaMA3-8B require 20.85 and 19.3, respectively. However, LLaMA-70B models took much longer, requiring 273.5-320.2 hours. Dataset characteristics such as note length impacted the training time; for example, MIMIC-III (2,106 tokens) took significantly more time per note than MTSamples (603 tokens). BiomedBERT outperformed all LLaMA models in inference, processing notes 5-28 times faster. Among LLaMA models, LLaMA-3-8B was quickest, while the 70B models were slowest. These findings highlight a trade-off: larger LLaMA models offer higher accuracy but are slower, while BiomedBERT is more efficient for time-sensitive applications.
LLaMA2- 7B LLaMA2- 13B LLaMA2- 70B LLaMA3- 8B LLaMA3- 70B Biomed BERT
Total Training time (hours)
UTP 38.3 65.5 294.5 35.5 251.5 89.0
MTSamples 1.8 3.0 13.5 1.6 11.5 4.1
MIMIC-III 1.6 2.7 12.3 1.5 10.5 3.7
Training GPU hours per epoch (hours)
UTP 19.2 32.8 147.3 17.8 125.8 4.5
MTSamples 0.9 1.5 6.8 0.8 5.8 0.2
MIMIC-III 0.8 1.4 6.2 0.8 5.3 0.2
Training time per note (seconds/note)
UTP 106.9 182.4 820.5 98.9 700.8 248.0
MTSamples 65.8 112.3 505.2 60.9 431.5 152.7
MIMIC-III 229.9 392.6 1,765.4 212.8 1,507.9 533.7
Total Inference time (seconds)
UTP 725.7 1,271.4 3,009.2 582.1 2,980.9 145.8
MTSamples 266.0 468.5 1,157.9 232.4 1,142.0 22.3
MIMIC-III 471.0 815.8 1,998.6 406.9 2,013.9 47.2
i2b2 398.8 695.4 1,753.5 335.4 1,786.2 81.6
Inference timer per note (seconds/note)
UTP 14.5 25.4 60.2 11.6 59.6 2.9
MTSamples 5.3 9.4 23.2 4.6 22.8 0.4
MIMIC-III 18.8 32.6 79.9 16.3 80.6 1.9
i2b2 8.0 13.9 35.1 6.7 35.7 1.6
  • Average training note lengths: UTP: 978 tokens/note; MTSamples: 603 tokens/note; MIMIC-III: 2106 tokens/note. Average test note lengths: UTP: 1,022 tokens/note; MTSamples: 655 tokens/note; MIMIC-III: 2,086 tokens/note; i2b2: 822 tokens/note. *70B models need two A100 80GB GPUs for inference.
Table 4: Training and test times, and processing throughputs across models and datasets.
3.4 Error analysis
Automatic and manual error analyses on LLaMA and BiomedBERT outputs across four datasets revealed that LLaMA-70B models have fewer False Negatives (FN) but slightly higher False Positives (FP) than smaller models, indicating a precision-recall trade-off with model size. BiomedBERT, while showing higher FN and FP rates, has fewer boundary errors, suggesting better accuracy in detecting text boundaries. Type errors were consistently low for all models, confirming effective classification once detected. Manual analysis of 40 sentences per dataset found a 7.5% annotation error rate, where models identified entities missed by human annotators, indicating their potential to improve pre-annotation processes.
3.5 Package Implementation
To support real-world application, we created an open-source clinical IE tool, ”Kiwi,” featuring our LLaMA-based and BiomedBERT-based models. Kiwi, available at https://kiwi.clinicalnlp.org/, is Python-based and packaged as a Docker image for cross-platform compatibility. It includes a CPU version with BiomedBERT and a GPU version with both BiomedBERT and LLaMA models. Kiwi docker image supports batch processing, and an online demo with a GUI is also provided on the website for testing purposes (see Figure 2 for interface screenshots).
Figure 2: Screenshots of Kiwi’s Web Interface (A) and Command Line Interface (B).
4 Discussion
This study presents the first comprehensive evaluation of instruction-tuned LLMs for clinical IE tasks using diverse institutional datasets, comparing them with previous SOTA BERT model. We address several limitations in existing LLM-based clinical IE research: previous studies often focused on datasets with limited entity types and relation types 31; 32, used small corpora for evaluation 43, and primarily tested on similar types of clinical texts 29; 31. By developing a broad, multi-institutional corpus including a wide range of medical contexts, entity types, and relation types, we offer a more comprehensive and generalized assessment of LLMs’ capabilities in real-world clinical IE scenarios.
Generative AI, especially LLMs, have gained immense popularity and have opened up vast potential across various industries. In healthcare, there is growing interest in utilizing LLMs to develop and implement IE systems that support both clinical research and practice. However, our study highlights that several challenges still impede the broad adoption of LLMs for clinical IE tasks. Despite their impressive performance, issues such as high computational costs and slow processing throughputs remain significant barriers. In the following sections, we discuss how to choose between LLMs and traditional deep learning approaches for clinical IE applications, factoring in performance, throughput requirements, and the availability of computational resources.
Performance and Generalizability: The consistent outperformance of LLaMA models, especially on the low-resource datasets (MTSamples, and MIMIC-III) and on unseen dataset (i2b2), underscores their potential to address the critical challenge of variability in clinical documentation 44. This variability, stemming from factors such as local practices, specializations, and patient demographics, has hindered the widespread adoption of NLP solutions in healthcare. The ability of instruction-tuned LLaMA models to generalize effectively across these variations suggests a promising path towards more universally applicable clinical NLP tools.
However, it’s important to note that while the performance gains are significant, they are not uniform across all tasks and datasets. When abundant annotated data are available at a healthcare system (e.g., UTP), the performance improvements of LLaMA are modest when compared with the BiomedBERT model (0.932 vs. 0.921 in F1 for NER and 0.937 vs 0.898 in F1 for RE), suggesting that deep learning models like BERT remain competitive in this scenario. Considering the low cost and high throughput of BERT models, we would recommend BERT models over LLMs when substantial annotated training data is already available for the IE task. However, when annotated training data is limited to a healthcare system (e.g., the case of i2b2), LLMs could be considered first, as they show better generalizability, indicating less annotation is needed to build the IE system. This highlights the need for careful consideration of the specific use case and available data when selecting a model for deployment.
Throughput: Our study highlights a significant difference in processing throughput between LLM-based and BERT-based systems, which is crucial for real-world clinical IE applications. In settings where real-time processing is required, such as during patient care, slow NLP systems are impractical as users cannot wait for delayed results. Similarly, in research environments handling millions of clinical notes, the lower processing throughput of LLMs makes them less feasible, even if they offer slightly better performance. Ultimately, the choice between LLMs and BERT-based systems must consider both performance and throughput, as the latter is often task-specific and essential in many clinical and research scenarios.
Computing Resources: The comparative analysis of computational requirements reveals important considerations for both inference and training. For inference deployment scenarios: (1) In high-resource settings with multiple NVIDIA A100 GPUs or better, LLaMA-3-70B offers the best performance; (2) In balanced settings with a single A100 GPU or equivalent, LLaMA-3-8B provides an attractive middle ground, achieving significant performance improvements; (3) For resource-constrained settings, BiomedBERT remains viable, requiring less resources while maintaining competitive performance on in-domain datasets. For training considerations, institutions have several options: (1) Using parameter-efficient tuning on larger LLaMA models (70B) demands significant resources but achieves the best performance; (2) Training smaller LLaMA models (7B/8B/13B) offers a balanced approach with competitive results; (3) Traditional fine-tuning of BiomedBERT provides the most resource-efficient though with lower performance on cross-institutional data. For institutions with extensive computational resources, pre-training domain-specific models like Me-LLaMA represents another option, though this requires more, making it practical only for large-scale research environments 45. These trade-offs underscore the importance of carefully matching computational capabilities with specific use case requirements.
Another significant contribution of this study is to release the Kiwi package. To the best of our knowledge, Kiwi is the first open-source clinical IE system based on LLMs. Kiwi includes both our instruction-tuned LLaMA-based and BiomedBERT-based models, available in both CPU and GPU versions to accommodate the diverse computational resources of different healthcare institutions. By offering different versions of models, Kiwi enables institutions to select the most appropriate model based on their specific needs, available resources, and use cases according to our suggestions. As a pioneering LLM-based clinical IE system, Kiwi represents a significant milestone in the application of generative AI to clinical IE. By making our system publicly available, we aim to accelerate research in LLM-based clinical IE by providing researchers with a ready-to-use framework. This will enable the medical NLP community to use our models directly, customize models upon our work, and develop more advanced clinical IE solutions.
While our study provides valuable insights into the potential of LLMs for clinical IE, future research should explore techniques to distill the knowledge of large LLMs into more compact, resource-efficient models 46; 47. This could help bridge the gap between the performance of large LLMs and the resource constraints of many healthcare settings. Future work should also explore how continual training on institution-specific datasets could further improve performance. This approach could allow individual healthcare institutions to refine the models’ performance on their unique documentation styles and terminologies, potentially achieving even better results than those observed in our cross-institutional evaluation. Several limitations of our study should be noted. First, our current information extraction approach operates at the sentence level, requiring clinical notes to be segmented into individual sentences before processing. This approach may not fully leverage one of the key advantages of LLMs: their ability to handle longer sequences. Future research could explore IE at the section or document level, potentially capturing more complex cases such as inter-sentence relationships. Second, our study utilized LLaMA-2/3 as base models, which were not specifically pretrained on medical texts. Future work could investigate instruction-tuning models based on domain-specific foundational LLMs, such as Me-LLaMA 45, which may yield better performance due to their specialized medical knowledge.
Data availability
The MTSamples dataset can be directly accessed on our GitHub repository. The MIMIC-III and 2010 i2b2 datasets, we are providing entity offsets and texts without the original text files on our GitHub repository.
Code availability
Our annotation guidelines and code used in this study are available at: https://github.com/BIDS-Xu-Lab/Kiwi-LLaMA
Acknowledgments
This study is partially supported by the following grants: NIA RF1AG072799, R01AG080429, and R01AG078154.
Contributions
Y.H. and H.X. contributed to the conception and design of the study. Y.H. and X.Z. performed the experiments and analyzed the data. Y.Z. led the data annotation. Y.H., Q.C., X.Z., and H.X. contributed to the interpretation of the results. Y.H., V.J., and J.H. contributed to the development of the Kiwi package. Y.H., J.H., V.J., V.K., and R.L.W. contributed to the interface development. Y.H., Q.C., X.Z, X.J, K.R., V.K., and H.X. drafted the manuscript and revised it. All authors approved the final version of the manuscript and agreed to be accountable for all aspects of the work.
Supplementary Information:
Annotation process:
For our annotation process, we engaged five annotators to ensure reliable data labeling for entity and relation recognition tasks. Annotator A, a medical student with extensive annotation experience, supervised the entire annotation process and helped resolve inconsistencies, making them the benchmark for annotation quality.
During the training phase, average F1 scores for entity and relation recognition across datasets showed the following results:
Annotator A vs. B: 82.5%
Annotator A vs. C: 79.3%
Annotator A vs. D: 78.1%
Annotator A vs. E: 75.4%
For the large-scale annotation tasks, each annotator will be assigned approximately 25 files per round, with 1-3 files designated for review and error analysis by Annotator A to maintain quality control. Our annotation guidelines, are available in our GitHub repository https://github.com/BIDS-Xu-Lab/Kiwi-LLaMA.
{xltabular}
p0.15—X Task Prompt
NER ### Task
Your task is to generate an HTML version of an input text, using HTML <span >tags to mark up specific entities.
### Entity Markup Guides:
Use <span class=”problem”>to denote a medical problem.
Use <span class=”treatment”>to denote a treatment.
Use <span class=”test”>to denote a test.
Use <span class=”drug”>to denote a drug.
### Entity Definitions:
Medical Problem: The abnormal condition that happens physically or mentally to a patient.
Treatment: The procedures, interventions, and substances given to a patient for treating a problem.
Drug: Generic or brand name of a single medication or a collective name of a group of medication.
Test: A medical procedure performed (i) to detect or diagnose a problem, (ii) to monitor diseases, disease processes, and susceptibility, or (iii) to determine a course of treatment.
### Input Text: Ortho Micronor 0.35 MG …
### Output Text: <span class=”drug”>Ortho Micronor </span>0.35 MG …
RE for problem entities ### Task
Your task is to mark up modifier entities related to the entity marked with <span >tag in the input text.
### Entity Markup Guides:
Use <span class=”uncertain”>to denote a measure of doubt.
Use <span class=”condition”>to denote a phrase that indicates the problems existing in a certain situation.
Use <span class=”subject”>to denote the person entity who is experiencing the disorder.
Use <span class=”negation”>to denote the phrase that indicates the absence of an entity.
Use <span class=”bodyloc”>to denote the location on the body where the observation is present.
Use <span class=”severity”>to denote the degree of intensity of a clinical condition.
Use <span class=”temporal”>to denote a calendar date, time, or duration related to a problem.
Use <span class=”course”>to denote the development or alteration of a problem.
### Input Text: probable <span class=”problem”>left paravertebral dilated vascular structure </span>…
### Output Text: <span class=”uncertain”>probable </span>left <span class=”bodyloc”>paravertebral </span>dilated vascular structure …
RE for treatment entities ### Task
Your task is to mark up modifier entities related to the entity marked with <span >tag in the input text.
### Entity Markup Guides:
Use <span class=”temporal”>to denote a calendar date, time, or duration related to a treatment.
Use <span class=”negation”>to denote the phrase that indicates the absence of an entity.
### Input Text: No <span class=”treatment”>further intervention </span>was done .
### Output Text: <span class=” negation”>No </span>further intervention was done .
RE for test entities ### Task
Your task is to mark up modifier entities related to the entity marked with <span >tag in the input text.
### Entity Markup Guides:
Use <span class=”labvalue”>to denote a numeric value or a normal description of the result of a lab test.
Use <span class=”reference_range”>to denote the range or interval of values that are deemed as normal for a test in a healthy person.
Use <span class=”negation”>to denote the phrase that indicates the absence of an entity.
Use <span class=”temporal”>to denote a calendar date, time, or duration related to a test.
### Input Text: <span class=”test”>Hgb </span>10.6 gm / dL
### Output Text: Hgb <span class=”labvalue”>10.6 gm / dL </span>
RE for drug entities ### Task
Your task is to mark up modifier entities related to the entity marked with <span >tag in the input text.
### Entity Markup Guides:
Use <span class=”form”>to denote the form of drug.
Use <span class=”frequency”>to denote the frequency of taking a drug.
Use <span class=”dosage”>to denote the amount of active ingredient from the number of drugs prescribed.
Use <span class=”duration”>to denote the time period a patient should take a drug.
Use <span class=”strength”>to denote the amount of active ingredient in a given dosage form.
Use <span class=”route”>to denote the way by which a drug, fluid, poison, or other substance is taken into the body.
Use <span class=”negation”>to denote the phrase that indicates the absence of an entity.
Use <span class=”temporal”>to denote a calendar date, time, or duration related to a drug.
### Input Text: <span class=”drug”>Ortho Micronor </span>0.35 MG …
### Output Text: Ortho Micronor <span class=”strength”>0.35 MG </span>…
Complete instruction examples for NER and RE
{xltabular}
—l—l—*12c— Dataset Model NER   RE
Exact Match   Relax Match   Exact Match   Relax Match
P R F1 P R F1 P R F1 P R F1
UTP LLaMA-2-7b 0.929 0.930 0.929 0.962 0.964 0.963 0.911 0.922 0.916 0.926 0.937 0.931
LLaMA-2-13b 0.930 0.934 0.932 0.961 0.967 0.964 0.903 0.928 0.915 0.917 0.944 0.930
LLaMA-2-70b 0.927 0.934 0.931 0.960 0.967 0.964 0.902 0.933 0.918 0.917 0.949 0.933
LLaMA-3-8b 0.933 0.926 0.929 0.968 0.961 0.965 0.930 0.943 0.936 0.944 0.959 0.952
LLaMA-3-70b 0.928 0.936 0.932 0.960 0.969 0.964 0.925 0.949 0.937 0.940 0.965 0.953
BiomedBERT 0.918 0.924 0.921 0.953 0.962 0.957 0.853 0.949 0.898 0.866 0.967 0.914
MTSamples LLaMA-2-7b 0.854 0.867 0.860 0.916 0.931 0.923 0.787 0.783 0.785 0.865 0.862 0.864
LLaMA-2-13b 0.861 0.874 0.868 0.921 0.936 0.928 0.773 0.814 0.793 0.849 0.897 0.872
LLaMA-2-70b 0.863 0.879 0.871 0.919 0.938 0.928 0.770 0.823 0.795 0.848 0.907 0.877
LLaMA-3-8b 0.869 0.869 0.869 0.931 0.931 0.931 0.774 0.801 0.787 0.854 0.883 0.868
LLaMA-3-70b 0.868 0.885 0.876 0.924 0.944 0.934 0.771 0.820 0.795 0.857 0.910 0.883
BiomedBERT 0.830 0.836 0.833 0.904 0.915 0.910 0.570 0.815 0.670 0.634 0.920 0.751
MIMIC-III LLaMA-2-7b 0.841 0.835 0.838 0.931 0.921 0.926 0.892 0.764 0.823 0.951 0.815 0.878
LLaMA-2-13b 0.846 0.849 0.847 0.932 0.934 0.933 0.880 0.790 0.833 0.942 0.843 0.890
LLaMA-2-70b 0.843 0.850 0.847 0.929 0.936 0.933 0.884 0.819 0.850 0.943 0.872 0.906
LLaMA-3-8b 0.849 0.838 0.843 0.936 0.923 0.930 0.898 0.824 0.859 0.948 0.869 0.907
LLaMA-3-70b 0.854 0.856 0.855 0.938 0.940 0.939 0.901 0.820 0.858 0.955 0.866 0.908
BiomedBERT 0.814 0.805 0.810 0.919 0.903 0.911 0.777 0.842 0.808 0.823 0.900 0.860
i2b2 LLaMA-2-7b 0.844 0.848 0.846 0.922 0.920 0.921 0.783 0.683 0.730 0.873 0.749 0.806
LLaMA-2-13b 0.848 0.859 0.853 0.923 0.927 0.925 0.757 0.694 0.725 0.848 0.766 0.805
LLaMA-2-70b 0.854 0.867 0.860 0.922 0.930 0.926 0.749 0.723 0.736 0.839 0.798 0.818
LLaMA-3-8b 0.857 0.847 0.852 0.936 0.917 0.926 0.761 0.719 0.739 0.856 0.796 0.825
LLaMA-3-70b 0.870 0.874 0.872 0.932 0.931 0.932 0.761 0.727 0.744 0.853 0.804 0.828
BiomedBERT 0.791 0.805 0.798 0.887 0.905 0.896 0.649 0.766 0.703 0.730 0.849 0.785
Exact and relax performance by LLaMA-2, LLaMA-3, and BiomedBERT on NER and RE across four datasets.
{xltabular}
—l—l—*12c— Dataset Model NER   RE
Exact Match   Relax Match   Exact Match   Relax Match
P R F1 P R F1 P R F1 P R F1
UTP LLaMA-2-7b 0.781 0.743 0.761 0.875 0.852 0.863 0.820 0.650 0.725 0.890 0.704 0.786
BiomedBERT 0.656 0.666 0.661 0.786 0.828 0.806 0.538 0.704 0.610 0.571 0.773 0.657
MTSamples LLaMA-2-7b 0.792 0.825 0.808 0.879 0.921 0.900 0.751 0.711 0.730 0.846 0.798 0.821
BiomedBERT 0.796 0.812 0.804 0.872 0.906 0.889 0.499 0.741 0.596 0.566 0.853 0.680
MIMIC-III LLaMA-2-7b 0.816 0.821 0.819 0.915 0.920 0.918 0.880 0.780 0.827 0.937 0.832 0.882
BiomedBERT 0.819 0.817 0.818 0.912 0.923 0.917 0.724 0.786 0.754 0.775 0.848 0.810
i2b2 LLaMA-2-7b 0.786 0.799 0.793 0.888 0.901 0.894 0.744 0.643 0.690 0.858 0.721 0.783
BiomedBERT 0.761 0.806 0.783 0.847 0.918 0.881 0.589 0.686 0.634 0.692 0.793 0.739
Exact and relax performance by LLaMA-3-70B, and BiomedBERT on NER and RE across four datasets solely using MIMIC-III dataset for finetuning.